OpenStudy (anonymous):

which equation represents one of the steps of solving 6x^2-6x-12=0 by factoring? a. 6x+4=0 or x-3=o b.6x+2=0 or x-6=0 c.2x-2=0 or 3x+6=0 d.x-2=0 or x+1=0

OpenStudy (myname):

First you take out 6 6(x^2-x-2)=0 x^2-x-2=0 (x-2)(x+1)=0 So, it's d.x-2=0 0r x+1=0